# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@
Abstract. We propose and analyze a nonlinear age-structured multi-species model that serves as a unifying framework for ecological and biotechnological systems in complex environments (microbial communities, bioreactors, etc.). The formulation incorporates nonlocal intra- and interspecific interactions modulated by environmental covariates; under general assumptions on mortality, reproduction rates and interaction kernels, we establish existence, uniqueness and positivity of solutions. We illustrate the model’s practical relevance along two lines: (i) multi-species examples, notably a non-transitive (cyclic) competition model, for which we show that, under the model assumptions, a control applied to a single species can achieve global stabilization of the system; furthermore, verification of the Kalman condition in this context provides an essential theoretical prerequisite and highlights that this single control acts indirectly on all other species; and (ii) the population dynamics of malaria-vector mosquitoes, for which we develop two control strategies (biological and genetic) and, in the biological-control scenario, prove global asymptotic stability of the aquatic compartment by constructing an explicit Lyapunov function. Numerical simulations validate the theoretical results and compare the effectiveness of the proposed strategies in reducing vector density and malaria transmission.
Keywords : Predator-prey ; stabilization ; non-transitive competition ; backstepping control, biological-genetic control.
1. Introduction, motivation, assumptions and modeling
@@ 1.1. Introduction and motivation
To capture not only predation but also competition, parasitism, mutualism, and related interaction types, we consider model (1.4), which represents a multispecies trophic network with diverse interaction structures and intensities. Unlike most classical formulations, this model explicitly incorporates age (or stage) structure, thereby linking demographic processes to interspecific interactions. Age-structured approaches are crucial, as survival, fecundity, and vulnerability often vary strongly across life stages, juveniles typically suffer higher mortality and lower reproduction, while adults contribute disproportionately to recruitment.
@@ 1.2. Assumptions and Modeling
Hypothesis (H5) provides a general formulation of the nonlocal birth term in age‑structured models, without resorting to a specific explicit form, and encompasses the particular cases treated in [28, 41, 43, 45].
@@
1.3. Interpretation of the model
We now ask: What kinds of ecological interactions can system (1.4) capture? The nonlinearity $f$ encodes the functional roles of species and their pairwise (or higher‐order) interactions. In particular, (1.4) can model any combination of the canonical interaction types: Consumer–resource (prey–predator), Interspecific competition, Mutualism, Non‐transitive dynamics.
To bring the model closer to ecological reality, we also include:
  • Intraspecific density‐dependence. Even in the absence of predators, prey populations cannot grow without bound. We impose a logistic‐type regulation: as population size increases, fertility decreases and mortality increases, reflecting limited resources, waste accumulation, social stress, etc.
  • Specialist versus generalist predation. A specialist predator (feeding on one species) disappears if its sole prey goes extinct, whereas a generalist can persist on alternative resources. In the latter case, its abundance is effectively independent of any single prey species.
  • Pack hunting and group defense. Many predators (lions, hyenas, dholes) hunt cooperatively, while some prey (buffalo herds) employ collective defense. Thus, the per‐capita predation rate depends on the structure and size of both predator and prey subpopulations.
Concretely, by choosing different forms of $f$ (e.g. mass‐action, Holling‐type II or III, saturating responses), one recovers in particular : The classical Lotka–Volterra predator–prey model, Two predators sharing one prey, or one predator with two prey, Food‐chain and food‐web modules (including competition and mutualism), Models with prey defense mechanisms (refuge, alarm calls, herding).
@@ 1.3. Interpretation of the model
[…]
1.4. Main results
Stability result-general case (for $N+1$)
A global asymptotic stability result was established for a non-transitive age-structured competition model. This result was first established for the case of three species, then extended to four species, and, by induction, generalized to the case of $N+1$ species.
Stability result for mosquito control
In the context of malaria vector control, we established a global stability result for a non-autonomous logistic model. Since adult mosquitoes emerge from the aquatic stage of the mosquito population, this stability result was obtained through a reduced control strategy targeting exclusively the aquatic population.
Result on the well-posedness of the model
Prior to establishing the stability results, the existence and uniqueness of solutions for the general model were rigorously established for several variants of the nonlinear function $f$ and the nonlocal term. These results were derived using semigroup theory and the Banach fixed-point theorem.
1.5. Related works and novelty
  • Under suitable hypotheses on fertility, mortality, and the interaction function, we establish well‑posedness results for multiple system configurations. In this section, we have established a threshold for resilience and/or the maintenance of biodiversity. Under assumption (H5-i), we prove existence, uniqueness, continuity and positivity of solutions for $f\leq 0$ (Theorems 2.62.7). Moreover, if $f\geq 0$ but the initial datum satisfies
    (1.8) $Y_{0}(a)\;\geq\;\Gamma(a+t;a)\;>\;0\quad\text{a.e. }a\in(0,A),\;t\in(0,T),$
@@ 1.5. Related works and novelty
  • We establish a general result on the global asymptotic stability of multi-species non-transitive competition models using backstepping control. We show that a single control (reduced control) is sufficient to stabilize these systems. After proving global asymptotic stability for three species in non-transitive competition, then extending the result to four species and, more generally, to $N+1$ species, we introduce a reduced control (multiplicative, applied to a single species) that naturally influences the dynamics without altering the model’s fundamental structure. This control is constructed using fictitious controls, designed to progressively and recursively stabilize subsystems of the model until the global control $u$ is synthesized. Concretely, each fictitious control partially stabilizes a given state or subsystem. Our approach offers additional contributions beyond existing techniques and models, providing a robust theoretical framework for stabilizing multi-species non-transitive competition systems. Our approach also allows the construction of any number of non-transitive competing species without altering the model’s fundamental structure. A global stabilization analysis of more than two age-structured species, based on fictitious controls that partially and recursively stabilize Lyapunov functions, is novel in the literature and extends naturally to more general settings.
    The nature of the model and the methodology Carina Veil et al. [48], where control actions are simultaneously applied to both species (prey and predator). In their approach, stabilization is achieved through the introduction of an additional control term that specifically strengthens the predator dynamics. In their more recent work on a two-species competition model [49], stabilization via the backstepping method is obtained without resorting to the construction of fictitious controls, owing to the reduced number of species considered and probably also to the particular structure of the model.
@@
Now, let us mention some related works from the literature
Many studies employ Lyapunov functions to demonstrate the stability of predator–prey models. Yet the complexity of the equations often forces one to work with simplified (and sometimes unrealistic) versions of the model in order to construct these functions analytically. Despite these simplifications, these works shed light on the so‑called “species harvesting” technique, which guarantees the positivity of solutions and their global asymptotic stability around a nonzero trajectory. This approach offers promising avenues for the development of conservation strategies. Moreover, the issues of well‑posedness and stability analysis for an age‑structured predator–prey model are examined in detail in [9, 10], and spatial extensions appear in [46].
In the context of vector control, several age‐structured mosquito models incorporate biological control methods, such as the sterile insect technique and cytoplasmic incompatibility, to account for larval and adult stages. A limited number of works (e.g. [2]) also include predator–mosquito interactions, though it remains challenging to model a single predator that effectively regulates both aquatic and adult mosquito populations.
1.6. Organization
The structure of the paper is as follows. Section 2 is devoted to the well-posedness of the model. In particular, we establish well-posedness in three settings: the demographic case ($f\equiv 0$), the linear nonlocal case, and the nonlinear nonlocal case. A special instance is discussed in Remark 2.11, which further clarifies its relation to Theorems 2.6 and 2.7. Section 3, devoted to applications, addresses design and stabilization. After a general introduction to multi-species models, we propose a multi-species backstepping control strategy for a non-transitive competition model. We then present a control strategy for a four-compartment mosquito dynamics model (aquatic stages, juvenile females, mature females and wild males) including the release of genetically modified individuals. Numerical simulations illustrate the evolution of each population with respect to age and time. Section 4 is devoted to the conclusion and outlines directions for future research.
@@ 2. Mathematical Analysis
To cast system (1.4) into an abstract Cauchy problem, we work in the Hilbert space $\mathcal{H}_{2}$. Under hypotheses $(H_{1})$$(H_{4})$ and for any fixed initial datum $Y_{0}\in\mathcal{H}_{2}$ (or $\mathcal{H}_{1}$), define
@@
[…]
    It is well known (e.g., [19, 26]) that $\mathcal{A}_{i}$ is the infinitesimal generator of a strongly continuous semigroup on $L^{2}(0,A_{i}).$ Since each of its elements is an infinitesimal generator of a semigroup, $\mathcal{A}$ is itself an infinitesimal generator of a semigroup. ∎
For similar operators, the reader may refer to the examples in [17, 32].
We now introduce the adjoint operator $\mathcal{A}^{*}$ of $\mathcal{A}$. Define
@@ 2.1. Model without species interaction
where $Y_{0}\in\mathcal{H}_{2}$.
This is the classical Lotka–McKendrick (or Lotka–von Foerster) model for $N$ non‑interacting species. Although it omits ecological feedbacks, it is mathematically well‑posed and serves as the prototype for age‑structured dynamics (see [16, 23, 21, 17, 22]). In particular, Inaba [17, Proposition 2.4] introduced exactly this “multi‑state” version as a stable‑population process.
@@
[…]
    The operator $\mathcal{A}$ is the infinitesimal generator of a $C_{0}$ semigroup of contraction on $\mathcal{H}_{2}$ [17]. Then, from [13, Theorem I], [14] the proof is immediate. ∎
@@
[…]
    As demonstrated in [32, Theorem 4], [34, Theorem 4.3], the same technique can be effectively utilized to establish this proposition. ∎
@@
The constants $M$ and $w$ arise from the estimation of the semi group in [14, Lemma p.19], [17, Lemma B].
@@ 2.2. Case of a linear nonlocal term
In this section, only the function $f$ is nonlinear and the renewal equation satisfies the condition (H5-i). The following result assures the existence and uniqueness of mild solutions of (1.4) for Lipschitz continuous functions $f.$
@@
    (2.9) $(FY)(t)=\mathcal{T}(t)\bigl{(}y_{01}(.),\dots,y_{0N}(.)\bigr{)}^{T}+\displaystyle\int_{0}^{t}\mathcal{T}(t-s)\lfloor-f(Y(.,s))\rfloor ds,\qquad t\in(0,T),$
@@
    (2.10) $(FY_{1})(t)-(FY_{2})(t)=\displaystyle\int_{0}^{t}\mathcal{T}(t-s)\lfloor-f(Y_{1}(.,s))+f(Y_{2}(.,s))\rfloor ds,\qquad t\in(0,T),$
    and thanks to $(H_{3}),$
@@
    (2.13) $\|Y(t)-\bar{Y}(t)\|_{\mathcal{H}_{2}}\leq Me^{wt}\|Y_{0}-\bar{Y}_{0}\|_{\mathcal{H}_{2}}+MLe^{wt}\displaystyle\int_{0}^{t}e^{-ws}\|Y(s)-\bar{Y}(s)\|_{\mathcal{H}_{2}}ds\quad\text{for every}\quad t\in(0,T)$
@@
    The uniqueness and continuity follows from (2.14). ∎
@@
[…]
    Let us fixed $\bar{Y}\in\mathcal{H}_{1}$ and define the mapping $\delta$ such that $\delta(\bar{Y})=Y(a,t).$ Consider in $\mathcal{X}=C((0;T);\mathcal{H}_{1})$ the norm $\|\bar{Y}\|_{\mathcal{X}}=\sup_{t\in(0,T)}e^{-\lambda t}\|\bar{Y}\|_{\mathcal{H}_{1}}$ for any $\bar{Y}(.,t)\in\mathcal{H}_{1}$, which is equivalent to the usual norm in $\mathcal{H}_{1}$ with $\lambda,$ a positive constant that will be made precise later.
@@
    and hypothesis $(H_{3})$ allows to obtain this estimate

### v2

# Stabilization by Harvesting in Age-Structured Trophic Networks

@@
Abstract. We propose and analyze a nonlinear age-structured multi-species model that serves as a unifying framework for ecological and biotechnological systems in complex environments (microbial communities, bioreactors, etc.). The formulation incorporates nonlocal intra- and interspecific interactions modulated by environmental covariates; under general assumptions on mortality, reproduction rates and interaction kernels, we establish existence, uniqueness and positivity of solutions. We illustrate the model’s practical relevance along two lines: multi-species examples, notably a non-transitive (cyclic) competition model, for which we show that, under the model assumptions, a control applied to a single species can achieve global stabilization of the system; and the population dynamics of malaria-vector mosquitoes, for which we develop two control strategies (biological and genetic) and, in the biological-control scenario, prove global asymptotic stability of the aquatic compartment by constructing an explicit Lyapunov function. Numerical simulations validate the theoretical results and compare the effectiveness of the proposed strategies in reducing vector density and malaria transmission.
Keywords : Predator-prey ; non-transitive competition ; stabilization ; backstepping control, biological-genetic control.
1. Introduction
@@ 1.1. Introduction and motivation
To capture not only predation but also competition, parasitism, mutualism, and related interaction types, we consider model (1.4), which represents a multispecies trophic network with diverse interaction structures and intensities. Unlike most classical formulations, this model explicitly incorporates age (or stage) structure, thereby linking demographic processes to interspecific interactions. Age-structured approaches are crucial, as survival, fecundity, and vulnerability often vary strongly across life stages, juveniles typically suffer higher mortality and lower reproduction, while adults contribute disproportionately to recruitment. For further details, the reader is invited to consult [35] as an example.
@@ 1.2. Problem setting
Assumption (H5) provides a general formulation of the non-local birth term in age‑structured models, without resorting to a specific explicit form, and includes the particular cases treated in [29, 44, 46, 47].
@@
Interpretation of the model
System (1.4) captures the full spectrum of canonical biotic interactions by encoding species’ roles and pairwise (or higher-order) effects in the nonlinearity $f$: consumer–resource (prey–predator), interspecific competition, mutualism and non-transitive dynamics. To increase realism the model also includes intraspecific density-dependence (logistic-type regulation of fertility and mortality to prevent unbounded growth), specialist versus generalist predation (specialists collapse with prey loss; generalists persist on alternative resources), and collective effects such as pack hunting and group defence (per-capita predation rates depend on predator and prey group structure and size). By choosing appropriate functional forms for $f$ (mass-action, Holling II-III, saturating responses, etc.) one recovers standard modules: Lotka–Volterra pairs, two predators-one prey or one predator-two prey configurations, food-chain and food-web motifs (including competition and mutualism), and models incorporating prey defence.
@@
The age-structured multi-population framework, via appropriate choices of $D(a)$ and $f$, provides a unified and flexible formalism to model and analyze stability, persistence, and control strategies in systems where demographic structure and non-local interactions are crucial. Although (1.4) was originally devised for animal ecology, it admits a broader interpretation when the notions of “birth,” “death,” and “interaction” are redefined: “species” can then represent agents, firms, or states, and the same framework captures complex dynamics in epidemiology, forest ecosystems, economics, geopolitics, or chemistry [37, 40, 43, 45, 47].
Main results
In this section we briefly describe our strategy and the principal results :
Global stability result - general $N$-species case
We establish a global asymptotic stability result for a class of non-transitive age-structured competition models. We first establish the result for three and four species, and then generalize it, by construction, to any $N$-species system via a constructive backstepping design.
Global stabilization result – mosquito control
We prove global stabilization of a non-autonomous logistic mosquito population model using a reduced control strategy acting exclusively on the aquatic stage.
Related works and novelty
Let us mention some related works from the literature
Many studies employ Lyapunov functions to demonstrate the stability of predator–prey models. Yet the complexity of the equations often forces one to work with simplified (and sometimes unrealistic) versions of the model in order to construct these functions analytically. Despite these simplifications, these works shed light on the so‑called “species harvesting” technique, which guarantees the positivity of solutions and their global asymptotic stability around a nonzero trajectory. This approach offers promising avenues for the development of conservation strategies. Structured models of competitive or predator-prey type have been extensively studied in the literature; for example, see the recent works of Carina Veil et al. [49, 50]. In [49], stabilization of the two-species predator–prey model is achieved by introducing an additional control term that specifically strengthens the predator dynamics. In [50], which addresses a two-species competition model, stabilization via the backstepping method is obtained without recourse to fictitious controls, owing to the small number of species considered and the particular structure of the model. Moreover, the issues of well‑posedness and stability analysis for an age‑structured predator–prey model are examined in detail in [9, 10].
In addition to work on age-structured predator–prey systems, the literature includes contributions on single-species age-structured models, focusing on the analysis of asymptotic stability and the design of robust control laws, as in [15, 27, 28].
In the context of vector control, several age‐structured mosquito models incorporate biological control methods, such as the sterile insect technique and cytoplasmic incompatibility, to account for larval and adult stages [6, 11, 30, 31, 41]. A limited number of works (e.g. [2]) also include predator–mosquito interactions, though it remains challenging to model a single predator that effectively regulates both aquatic and adult mosquito populations.
Contributions
  • Under suitable assumptions on fertility, mortality, and the interaction function, we establish well‑posedness results for multiple system configurations. In this section, we have established a threshold for resilience and/or the maintenance of biodiversity. Under assumption (H5-i), we prove existence, uniqueness, continuity and positivity of solutions for $f\leq 0$ (Theorems 2.62.7). Moreover, if $f\geq 0$ but the initial datum satisfies
    (1.8) $Y_{0}(a)\;\geq\;\Gamma(a+t;a)\;>\;0\quad\text{a.e. }(a,t)\in(0,A)\times(0,T),$
@@
  • Non-transitive competition, analogous to the “rock–paper–scissors” game, refers to a cyclic interaction network in which no species is globally dominant. Such dynamics tend to promote cyclic coexistence and complicate the natural stabilization of the system. Our work demonstrates that a single multiplicative control applied to one species is sufficient to globally stabilize the multi-species model under non-transitive competition. This control acts directly on the targeted species and indirectly influences the others through the cyclic network, ensuring convergence of the entire system toward a stable equilibrium.
    The methodology relies on a constructive design: fictitious controls are successively defined to partially stabilize subsystems of the model until the final global control is synthesized. Each fictitious control targets a specific state or subsystem while preserving the cyclic structure of the system. This procedure allows for the explicit construction of a common Lyapunov function, guaranteeing the asymptotic convergence of all species while preserving essential biological properties (positivity and boundedness of solutions).
    In this work, we go beyond the two-species framework to address multi-species stabilization, focusing specifically on non-transitive competition models with $N$ species. Existing methods for two species exploit low dimensionality and particular structural features that do not generalize easily. Increasing the number of species and introducing non-transitive interactions lead to cross-couplings and nonlinear terms, whose control requires a design that exploits the topological properties of the model and careful estimation of the interaction terms.
    The novelty of this work lies in several aspects:
    1. *
      Single control for $N$ species: we show that a multiplicative control applied to one species ensures the global stabilization of the system for any number of non-transitive species. The validity is first established for three and then four species, and extended to the general case through a constructive backstepping design.
    2. *
      Incremental backstepping procedure: the design relies on fictitious controls that gradually stabilize each subsystem while preserving the model’s topology. This backstepping scheme leads to the stepwise synthesis of a global Lyapunov function guaranteeing asymptotic stability.
    3. *
      This strategy is novel in the literature. Indeed, our approach preserves the non-transitive structure of species interactions: the applied control does not alter the cyclic dominance relations, while it qualitatively modifies the system dynamics so that trajectories converge to a globally stable equilibrium. Unlike existing studies that focus either on sustaining cycles or on directly stabilizing an equilibrium, our method achieves both simultaneously. Moreover, it guarantees positivity and boundedness of populations, thereby respecting ecological constraints and preserving the model’s cyclic topology.
    4. *
      Novel extension beyond two species: to our knowledge, very few results in the literature address the global stabilization of an arbitrary number of age-structured non-transitive species. Our framework fills this gap by providing a minimalist and cost-effective solution (a single control applied to one species, sufficient to stabilize the entire system) suitable for biological or management applications.
    The combination of these elements yields a robust global stabilization strategy, ensuring convergence of any multi-species non-transitive system to equilibrium without compromising its cyclic dynamics. This innovative backstepping approach offers a minimalist and feasible control, opening new perspectives for the regulation of complex biological communities.
@@
Organization
Our work is organized as follows. Section 2 is devoted to the well-posedness of the model. In particular, we establish well-posedness in three settings: the demographic case ($f\equiv 0$), the linear nonlocal case, and the nonlinear nonlocal case. A special instance is discussed in Remark 2.11, which further clarifies its relation to Theorems 2.6 and 2.7. Section 3, devoted to applications, addresses design and stabilization. After a general introduction to multi-species models, we propose a multi-species backstepping control strategy for a non-transitive competition model. We then present a control strategy for a four-compartment mosquito dynamics model (aquatic stages, juvenile females, mature females and wild males) including the release of genetically modified individuals. Numerical simulations illustrate the evolution of each population with respect to age and time. Section 4 is devoted to the conclusion and outlines directions for future research.
@@ 2. Mathematical Analysis
To cast system (1.4) into an abstract Cauchy problem, we work in the Hilbert space $\mathcal{H}_{2}$. Under hypotheses (H1)(H1)$(H4)$ and for any fixed initial datum $Y_{0}\in\mathcal{H}_{2}$ (or $\mathcal{H}_{1}$), define
@@
  • Proof of Lemma 2.1 :
    It is well known (e.g., [20, 26]) that $\mathcal{A}_{i}$ is the infinitesimal generator of a strongly continuous semigroup on $L^{2}(0,A_{i}).$ Since each of its elements is an infinitesimal generator of a semigroup, $\mathcal{A}$ is itself an infinitesimal generator of a semigroup.
For similar operators, the reader may refer to the examples in [18, 33]. We now introduce the adjoint operator $\mathcal{A}^{*}$ of $\mathcal{A}$. Define
@@ 2.1. Model without species interaction
where $Y_{0}\in\mathcal{H}_{2}$. This is the classical Lotka–McKendrick (or Lotka–von Foerster) model for $N$ non‑interacting species. Although it omits ecological feedbacks, it is mathematically well‑posed and serves as the prototype for age‑structured dynamics (see [17, 18, 23, 27, 22]). In particular, Inaba [18, Proposition 2.4] introduced exactly this “multi‑state” version as a stable‑population process.
@@
  • Proof of Proposition 2.3 :
    The operator $\mathcal{A}$ is the infinitesimal generator of a $C_{0}$ semigroup of contraction on $\mathcal{H}_{2}$ [18]. Then, from [13, Theorem I], [14] the proof is immediate.
@@
  • Proof of Proposition 2.4 :
    As demonstrated in [33, Theorem 4], [36, Theorem 4.3], the same technique can be effectively utilized to establish this proposition.
@@
The constants $M$ and $w$ arise from the estimation of the semi group in [14, Lemma p.19], [18, Lemma B].
@@ 2.2. Case of a linear nonlocal term
In this section, only the function $f$ is nonlinear and the renewal equation satisfies the condition (H5-i). The following result assures the existence and uniqueness of mild solutions of (1.4) for Lipschitz continuous functions $f.$ An example of a typical age-structured model can be found in [33].
@@
    (2.9) $(FY)(t)=\mathcal{T}(t)\bigl(y_{01}(.),\dots,y_{0N}(.)\bigr)^{T}+\displaystyle\int_{0}^{t}\mathcal{T}(t-s)\lfloor-f(Y(.,s))\rfloor ds,\hskip 18.49988ptt\in(0,T),$
@@
    (2.10) $(FY_{1})(t)-(FY_{2})(t)=\displaystyle\int_{0}^{t}\mathcal{T}(t-s)\lfloor-f(Y_{1}(.,s))+f(Y_{2}(.,s))\rfloor ds,\hskip 18.49988ptt\in(0,T),$
    and thanks to $(H3),$
@@
    (2.13) $\|Y(t)-\bar{Y}(t)\|_{\mathcal{H}_{2}}\leq Me^{wt}\|Y_{0}-\bar{Y}_{0}\|_{\mathcal{H}_{2}}+MLe^{wt}\displaystyle\int_{0}^{t}e^{-ws}\|Y(s)-\bar{Y}(s)\|_{\mathcal{H}_{2}}ds\qquad\text{for every}\qquad t\in(0,T)$
@@
    The uniqueness and continuity follows from (2.14).
@@
  • Proof of Theorem 2.7 :
    Let us fixed $\bar{Y}\in\mathcal{H}_{1}$ and define the mapping $\delta$ such that $\delta(\bar{Y})=Y(a,t).$ Consider in $\mathcal{X}=C((0;T);\mathcal{H}_{1})$ the norm
    $\|\bar{Y}\|_{\mathcal{X}}=\sup_{t\in(0,T)}e^{-\lambda t}\|\bar{Y}\|_{\mathcal{H}_{1}},\qquad\;\text{for any}\;\bar{Y}(.,t)\in\mathcal{H}_{1},$
    which is equivalent to the usual norm in $\mathcal{H}_{1}$ with $\lambda,$ a positive constant that will be made precise later.
@@
    and assumption $(H3)$ allows to obtain this estimate